OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10) a. y =x b.y=x+6 c.y=x-6 d.y=x+6 which one is it? a,b,c,or d?

OpenStudy (anonymous):

Do you want to know the option or the long answer?

OpenStudy (anonymous):

i want to know the option...is it a,b,c,or d?

OpenStudy (anonymous):

Sorry. I don't help that way.

OpenStudy (anonymous):

can you tell me anything?

OpenStudy (anonymous):

Yes, I can tell you many things. Giving you the option isn't one of them

OpenStudy (anonymous):

can you tell me what to do to get the answer...

OpenStudy (anonymous):

Yes, absolutely. the equation of a line in the slope-intercept form is y = mx +b, where m is the slope and b is the offset. It is given that the line is parallel to y = x + 8. This means that the slope of both lines is equal. In y = x + 8, can you tell me what is the slope, according to the slope-intercept form?

OpenStudy (anonymous):

4,10?

OpenStudy (anonymous):

no. y = mx + b ------> equation 1 y = x + 8 ------> equation 2 can you tell me if you see any similarities in equations 1 and 2?

OpenStudy (anonymous):

do i have to ad 8 to both sides?

OpenStudy (anonymous):

No no. Just tell me if you see any similarities.

OpenStudy (anonymous):

y-0x+8?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

0 times anything is 0.

OpenStudy (anonymous):

i see similarities ....i think their similar

OpenStudy (anonymous):

okay great. y = mx +b y = x +8. Can you see that m is 1 in the second equation and b is 8? Anyway, the two lines are parallel. This means that the second line has slope 1. this means, its equation in slope-intercept form is y = x + b. This line passes through the point (4,10) substitute x = 4 and y = 10 in the equation above and find b.

OpenStudy (anonymous):

ok ill do that now

OpenStudy (anonymous):

b=6 !?

OpenStudy (anonymous):

yes, very good! You are almost there! So, what is the correct option?

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

??

OpenStudy (anonymous):

right. very good!

OpenStudy (anonymous):

although b and d are the same equation according to your question.

OpenStudy (anonymous):

thank you soo much now i wont struggle with these types of questions

OpenStudy (anonymous):

you are welcome :)